What is the speed of a wavelength 354 nm with 2.55 * 10^19 hz traveling inside a vacuum?
Answer #1

Given:

Wavelength = 354 nm

= 354*10^-9 m

= 3.54*10^-7 m

f = 2.55*10^19 Hz

= 2.55*10^19 s-1

Use:

Speed = wavelength * frequency

= 3.54*10^-7 m * 2.55*10^19 s-1

= 9.03*10^12 m/s

Answer: 9.03*10^12 m/s